# Multiply Robust Estimation for Local Distribution Shifts with Multiple Domains

### v1

@@ B.2 Supporting Lemmas
[…]
For simplicity, we derive this result which is not conditional on a segment $s$, however, this can immediately be applied to the conditional case.
Proof.
[…]
where $\sigma_{i}$ are independent uniform {±$\{\pm 1\}$ random variables.
[…]
Lemma B.3.
Let $\widehat{w}_{s}$ be an estimator of the true importance weights. Then
$|R[f|Y,X,s,\widehat{w}_{s}]-R[f|Y,X,s,w_{s}]|\leq L\frac{\lVert\widehat{w}_{s}-w_{s}\rVert_{1}}{n_{s}}$
Proof.
The proof is a straightforward application of Holder’s inequality.
$|R[f|Y,X,s,\widehat{w}_{s}]-R[f|Y,X,s,w_{s}]|$ $=|\frac{1}{n_{s}}\sum_{i\in N_{s}}\ell(Y_{i},f(X_{i}))(\widehat{w}_{s}(X_{i},Y_{i})-w_{s}(X_{i},Y_{i}))|$
$\leq\text{max}_{i\in N_{s}}|\ell(Y_{i},f(X_{i}))|\frac{\lVert\widehat{w}_{s}-w_{s}\rVert_{1}}{n_{s}}$
$\leq L\frac{\lVert\widehat{w}_{s}-w_{s}\rVert_{1}}{n_{s}}$
B.4 Bounds on estimating the importance weights
Our upper bound relies on estimating the importance weights $\lVert\widehat{w}_{s}-w_{s}\rVert_{1}$.
We focus on two results, one for covariate shift and one for label shift, which allow for the estimation of each of these weights.
Lemma B.4 (Lemma 8 of (Reddi et al., 2015)).
Suppose that $\widehat{w}_{s}$ is an estimator for $w_{s}$ obtained by Penalized Risk Minimization (as in (Nguyen et al., 2010)). Suppose $\gamma_{n}=cn_{s}^{-2/(2+\tau)}$ for some $\tau>0$. Then with probability at least $1-\delta$
@@ B.4 Bounds on estimating the importance weights
Let $\mathbf{C}$ denote the expected confusion matrix computed under the train distribution $\mathbf{C}_{i,j}=P(\widehat{y}=i,y=j)$ where $\mathbf{C}$ is invertible where $\sigma_{min}$ it’s smallest eigenvalue. Then there is a universal constant $C$ such that for all $n>80\log(n)\sigma_{min}^{-2}$. Then with probability at least $1-\frac{5K}{n^{10}}$.
@@ C.1 Data splitting and cross validation
The primary reason for data splitting is when training the base models $\{h_{m}\}_{m=1}^{M}$ and fitting the stage 1 estimator using the same data may lead to overfitting. Alternatively, one may split the training data and preserve a holdout set (e.g. 20%) which will be used for estimating $\beta$. Then use the entirety of a segment’s training data to tune the estimate $\widehat{f}^{(s)}$. We provide a visualization in Figure 4. This splitting procedure can be simply generalized to a K-fold versions, as well as grouped versions for data types that are clustered by observational unit. Cross-validation details for the numerical experiments are included in Section C.2 and Section C.3
Figure 4: Data splitting for the multiply-robust estimator. We use a holdout set for refining the linear combination, and train the second stage model on the whole training fold segment.
@@ C.4 Learned Clusters by Segment
After learning computing a distance matrix using Algorithm 2 we use the ward linkage and pick the threshold that maximizes the number of base models $M$ before isolating a single segment.
UCI Datasets
@@ C.5 Model Performance Across Segments
Table 15: User City Dataset. Relative Cross Entropy to the Prduction model presented for each of the country segment.
Country XGB DR DR-SF MR MFSAN DADIL
Country 1 0.944 (0.222) 0.898 (0.217) 0.887 (0.218) 1.362 (0.475) 1.139 (0.253) 1.27 (0.283)
Country 2 0.921 (0.138) 0.876 (0.129) 0.89 (0.128) 0.847 (0.142) 1.008 (0.142) 1.635 (0.203)
Country 3 0.957 (0.035) 0.911 (0.032) 0.835 (0.035) 0.763 (0.035) 1.246 (0.038) 2.168 (0.056)
Country 4 0.951 (0.128) 0.912 (0.122) 0.891 (0.121) 0.826 (0.108) 1.101 (0.123) 1.835 (0.167)
Country 5 0.95 (0.258) 0.953 (0.259) 0.913 (0.246) 0.958 (0.271) 0.876 (0.242) 1.727 (0.424)
Country 6 0.962 (0.091) 0.91 (0.084) 0.937 (0.085) 0.864 (0.097) 0.995 (0.086) 2.135 (0.177)
Country 7 0.923 (0.068) 0.92 (0.068) 0.908 (0.069) 0.832 (0.073) 1.067 (0.074) 1.486 (0.104)
Country 8 0.975 (0.085) 0.924 (0.085) 0.931 (0.083) 0.911 (0.105) 1.042 (0.088) 2.353 (0.192)
Country 9 0.988 (0.109) 0.951 (0.102) 0.92 (0.098) 0.908 (0.103) 1.062 (0.104) 2.035 (0.164)
Country 10 0.996 (0.056) 0.946 (0.053) 0.966 (0.052) 0.916 (0.058) 1.234 (0.072) 1.524 (0.097)
Country 11 0.995 (0.035) 0.954 (0.034) 0.966 (0.032) 0.914 (0.034) 1.148 (0.04) 1.908 (0.072)
Country 12 0.965 (0.072) 0.932 (0.071) 0.942 (0.07) 0.865 (0.075) 1.051 (0.076) 1.972 (0.133)
Country 13 0.995 (0.052) 0.963 (0.05) 0.757 (0.061) 0.674 (0.065) 1.211 (0.059) 3.324 (0.133)
Country 14 1.004 (0.147) 0.994 (0.134) 0.882 (0.139) 0.978 (0.268) 2.016 (0.215) 2.243 (0.196)
Country 15 1.051 (0.472) 1.099 (0.465) 1.021 (0.416) 0.86 (0.322) 1.705 (0.551) 2.938 (1.209)
Country 16 1.048 (0.168) 0.963 (0.15) 0.868 (0.145) 0.807 (0.147) 1.071 (0.145) 1.448 (0.191)
Country 17 0.9 (0.192) 0.937 (0.207) 0.962 (0.204) 0.953 (0.237) 1.116 (0.215) 2.007 (0.412)
Country 18 1.024 (0.193) 0.984 (0.177) 0.984 (0.177) 0.956 (0.21) 1.129 (0.204) 1.531 (0.227)
Country 19 0.96 (0.054) 0.915 (0.051) 0.882 (0.05) 0.797 (0.053) 1.277 (0.062) 1.384 (0.071)
Country 20 0.97 (0.039) 0.932 (0.038) 0.934 (0.036) 0.864 (0.035) 1.153 (0.042) 1.845 (0.063)
Other 0.975 (0.02) 0.926 (0.019) 0.934 (0.018) 0.859 (0.019) 1.683 (0.027) 1.781 (0.033)

### v2

@@ B.4 Bounds on estimating the importance weights
Suppose that $\widehat{w}_{s}$ is an estimator for $w_{s}$ obtained by Penalized Risk Minimization (as in (Nguyen et al., 2010)). Suppose $\gamma_{n}=cn_{s}^{-2/(2+\tau)}$ for some $\tau>0$. Let $\eta:w_{s}\leq\eta$ denote the bound of the magnitude of the true importance weights. Then with probability at least $1-\delta$
@@
Let $\mathbf{C}$ denote the expected confusion matrix computed under the training distribution $\mathbf{C}_{i,j}=P(\widehat{Y}=i,Y=j)$ where $\mathbf{C}$ is invertible where $\sigma_{min}$ it’s smallest eigenvalue. Then there is a universal constant $C$ such that for all $n>80\log(n)\sigma_{min}^{-2}$. Then with probability at least $1-\frac{5K}{n^{10}}$.
@@
$\leq\frac{\sqrt{C}}{\sigma_{min}\left(\sum_{y}\tilde{w}^{2}(y)+K\right)}\sqrt{\frac{\log(n)}{n}}$
@@ C.1 Data splitting and cross validation
The primary reason for data splitting is to avoid overfitting that can occur when the same data is used for training the base models ${h_{m}}_{m=1}^{M}$ and subsequently fitting the stage 1 estimator. Alternatively, one may split the training data and preserve a holdout set (e.g. 20%) which will be used for estimating $\beta$, then use use the fine-tuning dataset to estimate $\widehat{f}^{(s)}$. We provide a visualization in Figure 4. This splitting procedure can be simply generalized to a K-fold versions, as well as grouped versions for data types that are clustered by observational unit (for example, the multiple candidate cities for each userid in our location dataset). Cross-validation details for the numerical experiments are included in Section C.2 and Section C.3.
Given a training and test dataset, we constructed a cross-validation procedure as follows:
  1. 1.
    Split the training data into $K$ folds.
  2. 2.
    For each $k$ in ${1,2,…K}$:
    1. (a)
      Use the $k^{th}$ fold as the validation fold, and combine the remaining folds to form the train fold.
    2. (b)
      Fit the MR estimator using algorithm 1 on the train fold, splitting it in a base train and a tune fold in the process.
    3. (c)
      Compute the importance weights on the validation fold to match the test data.
    4. (d)
      For each segment, evaluate the fine-tuned MR estimator on the weighted validation fold.
Figure 4: Data splitting for the multiply robust estimator. We use a holdout set for refining the linear combination, and train the second stage model on the whole training fold segment.
@@ C.4 Learned Clusters by Segment
After computing a distance matrix using Algorithm 2 we apply Ward’s linkage for hierarchical agglomerative clustering and select the threshold that maximizes the number of base models $M$ before isolating a single segment.
@@ C.5 Model Performance Across Segments
Table 15: User City Dataset. Relative Cross Entropy to the Production model presented for each of the country segment. We include MR - LC for our method which only includes the first stage estimators without the second stage fine-tuning.
Country XGB DR SF DR-SF MR MR - LC MFSAN DADIL
Country 1 0.944 (0.222) 0.898 (0.217) 0.961 (0.24) 0.887 (0.218) 1.362 (0.475) 1.550 (0.512) 1.139 (0.253) 1.270 (0.283)
Country 2 0.921 (0.138) 0.876 (0.129) 0.938 (0.134) 0.890 (0.128) 0.847 (0.142) 1.041 (0.153) 1.008 (0.142) 1.635 (0.203)
Country 3 0.957 (0.035) 0.911 (0.032) 0.855 (0.037) 0.835 (0.035) 0.763 (0.035) 0.842 (0.037) 1.246 (0.038) 2.168 (0.056)
Country 4 0.951 (0.128) 0.912 (0.122) 0.940 (0.131) 0.891 (0.121) 0.826 (0.108) 0.876 (0.111) 1.101 (0.123) 1.835 (0.167)
Country 5 0.950 (0.258) 0.953 (0.259) 0.919 (0.259) 0.913 (0.246) 0.958 (0.271) 1.120 (0.222) 0.876 (0.242) 1.727 (0.424)
Country 6 0.962 (0.091) 0.910 (0.084) 0.989 (0.092) 0.937 (0.085) 0.864 (0.097) 0.991 (0.099) 0.995 (0.086) 2.135 (0.177)
Country 7 0.923 (0.068) 0.920 (0.068) 0.956 (0.072) 0.908 (0.069) 0.832 (0.073) 0.943 (0.075) 1.067 (0.074) 1.486 (0.104)
Country 8 0.975 (0.085) 0.924 (0.085) 0.990 (0.085) 0.931 (0.083) 0.911 (0.105) 1.089 (0.107) 1.042 (0.088) 2.353 (0.192)
Country 9 0.988 (0.109) 0.951 (0.102) 0.949 (0.101) 0.920 (0.098) 0.908 (0.103) 1.013 (0.103) 1.062 (0.104) 2.035 (0.164)
Country 10 0.996 (0.056) 0.946 (0.053) 1.010 (0.056) 0.966 (0.052) 0.916 (0.058) 1.017 (0.056) 1.234 (0.072) 1.524 (0.097)
Country 11 0.995 (0.035) 0.954 (0.034) 0.995 (0.033) 0.966 (0.032) 0.914 (0.034) 1.042 (0.032) 1.148 (0.040) 1.908 (0.072)
Country 12 0.965 (0.072) 0.932 (0.071) 0.967 (0.072) 0.942 (0.07) 0.865 (0.075) 0.937 (0.067) 1.051 (0.076) 1.972 (0.133)
Country 13 0.995 (0.052) 0.963 (0.050) 0.763 (0.060) 0.757 (0.061) 0.674 (0.065) 0.751 (0.065) 1.211 (0.059) 3.324 (0.133)
Country 14 1.004 (0.147) 0.994 (0.134) 0.831 (0.153) 0.882 (0.139) 0.978 (0.268) 1.036 (0.273) 2.016 (0.215) 2.243 (0.196)
Country 15 1.051 (0.472) 1.099 (0.465) 1.177 (0.485) 1.021 (0.416) 0.86 (0.322) 0.956 (0.269) 1.705 (0.551) 2.938 (1.209)
Country 16 1.048 (0.168) 0.963 (0.150) 1.016 (0.179) 0.868 (0.145) 0.807 (0.147) 0.920 (0.141) 1.071 (0.145) 1.448 (0.191)
Country 17 0.900 (0.192) 0.937 (0.207) 0.948 (0.206) 0.962 (0.204) 0.953 (0.237) 1.013 (0.251) 1.116 (0.215) 2.007 (0.412)
Country 18 1.024 (0.193) 0.984 (0.177) 0.981 (0.173) 0.984 (0.177) 0.956 (0.21) 1.311 (0.284) 1.129 (0.204) 1.531 (0.227)
Country 19 0.960 (0.054) 0.915 (0.051) 0.917 (0.052) 0.882 (0.05) 0.797 (0.053) 0.874 (0.05) 1.277 (0.062) 1.384 (0.071)
Country 20 0.970 (0.039) 0.932 (0.038) 0.977 (0.037) 0.934 (0.036) 0.864 (0.035) 0.983 (0.035) 1.153 (0.042) 1.845 (0.063)
Other 0.975 (0.020) 0.926 (0.019) 0.990 (0.019) 0.934 (0.018) 0.859 (0.019) 0.972 (0.018) 1.683 (0.027) 1.781 (0.033)